suppose real GDP is growing at 4 percent, the money supply is growing at 11 percent, the velocity of money is constant, the real interest rate is 6 percent, what is the current inflation rate and nominal interest rate ? if the money supply growth rate is 15 percent, how will your answer in part (a) change? if your an investor, how would the change in the money supply growth affect your real profitability, assuming that you now receive the new nominal interest rate?
1
Expert's answer
2012-09-11T08:24:56-0400
a) M/P=Y/V
Y - growing by 4% so coeficient = 1.04
M - growing by 11% so coeficient = 1.11
V - constant =1
1.11/P=1.04/1
P=1.11/1.04≈1.067≈+6.7%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.06=(1+i)/(1+0.067)
i=(1.06*1.067)-1
i≈0.13102≈13.10%
______ ______ ______ ______

(b)
M/P=Y/V
Y - growing by 4% so coeficient = 1.04
M - growing by 15% so coeficient = 1.15
V - constant =1
1.15/P=1.04/1
P=1.15/1.04≈1.1057692≈+10.57%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.06=(1+i)/(1+0.1057)
i=(1.06*1.1057)-1
i≈0.1720≈17.20%
______ ______ ______ ______

i¹≈13.10%
i²≈17.20%
Δi=(17.20/13.10)-1≈
≈1.3129-1≈+31.29%
c)If money supply is growing real profitability is declining. For the same money Investor could afford less goods.
d) I would prefer fixed interest rate because it gives me understanding of stable earning money